# Constrained Portfolio Analysis in High Dimensions: Tracking Error and Weight Constraints

### v1

@@ 3 Tracking Error Constraint
with $\boldsymbol{m}\mathrel{\mathop{:}}p\times 1$ benchmark portfolio of weights, $\boldsymbol{w}$ is the portfolio of weights that is tracking the benchmark portfolio $\boldsymbol{m}$, $\boldsymbol{\mu}\mathrel{\mathop{:}}p\times 1$ vector of expected return on assets, and $TE$ is the tracking error. Alternatively we can write the optimization as
$\boldsymbol{w}_{d}^{*}\mathrel{\mathop{:}}=\max_{\boldsymbol{w}_{d}\in R^{p}}[\boldsymbol{\mu}^{\prime}\boldsymbol{w}_{d}-\frac{\Xi}{2}\boldsymbol{w}_{d}^{\prime}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{d}],\quad\mbox{subject to}\quad\boldsymbol{1}_{p}^{\prime}\boldsymbol{w}_{d}=0,$
@@
with $\boldsymbol{w}_{d}=\boldsymbol{w}-\boldsymbol{m}$, $\infty>\Xi>0$, where $\Xi$ can be roughly seen as a risk aversion parameter. The solution to above problem is, $\boldsymbol{\Theta}\mathrel{\mathop{:}}=\boldsymbol{\Sigma}_{y}^{-1}$, as in (1) of Bajeux-Besnainou et al. (2011)
@@
where $\frac{\boldsymbol{\Theta}\boldsymbol{\mu}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}}$ is the portfolio that maximizes Sharpe ratio, and $\frac{\boldsymbol{\Theta}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}$ is the standard global minimum variance portfolio. Let $c,C$ be positive constants. Then $\kappa$, represents the risk tolerance parameter, which is described in equation (5) of Bajeux-Besnainou et al. (2011), with range of $0<c\leq|\kappa|\leq C<\infty$. For risk averse investors $0<c\leq\kappa\leq C<\infty$. $\kappa,\boldsymbol{m}$ are given by investors and not estimated. The relation between $TE,\kappa,\Xi$ is given in Section A.2. The aim is to estimate optimal weights $\boldsymbol{w}_{d}^{*}$, and $\boldsymbol{w}^{*}\mathrel{\mathop{:}}=\boldsymbol{w}_{d}^{*}+\boldsymbol{m}$. In that respect we provide the following estimate
$\hat{\boldsymbol{w}}_{d}=\kappa\left[\frac{\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}}-\frac{\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}\boldsymbol{1}_{p}}\right],$
with $\hat{\boldsymbol{\mu}}=\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{y}_{t}$, with $y_{t},t=1,\cdots,T$ being the columns of $\boldsymbol{y}$ matrix. Note that since
@@
we have
$\|\hat{\boldsymbol{w}}_{d}-\boldsymbol{w}_{d}^{*}\|_{1}=\|\hat{\boldsymbol{w}}-\boldsymbol{w}^{*}\|_{1},$
given $\hat{\boldsymbol{w}}=\hat{\boldsymbol{w}}_{d}+\boldsymbol{m}$. We provide the following theorem, that shows that portfolio weights can be consistently estimated in a large portfolio with tracking error constraint.
Also we have two other metrics that we consider. Out-of-sample variance of a portfolio, and out-of-sample-Sharpe-Ratio corresponding to constrained portfolio weight estimates $\hat{\boldsymbol{w}}$. Specifically they are, the out of sample variance estimate of the constrained portfolio
@@
and the estimate of the out-of-sample-Sharpe-Ratio of the constrained portfolio
@@
The Sharpe ratio is $SR=\boldsymbol{w}^{*^{\prime}}\boldsymbol{\mu}/\sqrt{\boldsymbol{w}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}^{*}}$.
Define the rate $r_{w1}\mathrel{\mathop{:}}=\bar{s}^{2}l_{T}r_{T}K^{7/2}$. Note that $l_{T}\to 0$ and is defined in (2), and $r_{T}\to\infty$ with $T\to\infty$, and is defined in Assumption 6, and $\bar{s}$ is nondecreasing in $T$, and also $K$ is nondecreasing in $T$. We provide consistent estimation of portfolio weights and out-of-sample variance of the constrained portfolio.
Theorem 1.
@@
$\|\hat{\boldsymbol{w}}-\boldsymbol{w}^{*}\|_{1}=\|\hat{\boldsymbol{w}}_{d}-\boldsymbol{w}_{d}^{*}\|_{1}=O_{p}(r_{w1})=O_{p}(\bar{s}^{2}l_{T}r_{T}K^{7/2})=o_{p}(1).$
@@
$\left|\frac{\hat{\boldsymbol{w}}^{{}^{\prime}}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}}{\boldsymbol{w}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}^{*}}-1\right|=O_{p}(r_{w1}\bar{s}r_{T}K^{9/2})=O_{p}(\bar{s}^{3}l_{T}r_{T}^{2}K^{8})=o_{p}(1).$
@@
$\left|\left[\frac{\widehat{SR}}{SR}\right]^{2}-1\right|=O_{p}(r_{w1}\bar{s}r_{T}K^{9/2})=O_{p}(\bar{s}^{3}l_{T}r_{T}^{2}K^{8})=o_{p}(1).$
Remarks. 1. These are all new results in the literature and allows $p>T$, and both can grow to $\infty$ due to $l_{T}$ definition in (2). An example of a rate is given at the end of section 1. Corollary 3 of Caner et al. (2023) shows unconstrained maximum out of sample Sharpe ratio can be consistenly estimated with a rate of $\bar{s}l_{T}K^{2}$, so our Theorem 1(iii) here shows the effect of the tracking error constraint, and our estimation error is much larger.
@@
3. If the precision matrix of errors is non-sparse we can have $\bar{s}=p$. Then in this case $p<T$, and we need $p^{2}l_{T}r_{T}K^{7/2}=o(1)$, hence we need $p=o(T^{1/4})$, by using $l_{T}$ definition in (2) for Theorem 1(i). For Theorems 1(ii)-(iii), with non-sparse $\bar{s}=p$, we need $p=o(T^{1/6})$.
@@
4. We can also have an estimate for tracking error as $\widehat{TE}_{d}\mathrel{\mathop{:}}=\sqrt{\hat{\boldsymbol{w}}_{d}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{d}}$, as an out-sample estimate. It can easily be shown that $|(\frac{\widehat{TE}_{d}}{TE})^{2}-1|$ will converge in probability to zero by following the proof of Theorem 1(ii), and the rate will be the same as in Theorem 1(ii). This estimator is consistent.
@@
5. Note that estimating variance and Sharpe Ratio is more difficult compared to weight formation. This is clear by comparing Theorem 1(i) with Theorem 1(ii)(iii). We can also estimate in-sample Sharpe Ratio with our analysis as in Caner et al. (2023) but out-sample one is more related to portfolio advice in practice, hence we analyze out-sample here.
4 Joint Tracking Error and Weight Constraints
@@
In this section we analyze weight constraints for a portfolio. Let $R$ represent the indices of restrictions and we define the cardinality of the set $R$ as $r\mathrel{\mathop{:}}=|R|$. We allow for at least 1 restricted asset and also all the assets cannot be restricted. In other words, if the number of restrictions are $r$, $1\leq r<p$. To give an example, let $p=10$, and $R=\{1,5,7\}$, this means assets 1, 5, 7 are restricted in an 10 asset portfolio, and $r=3$. To make things a bit clear in notation, define $\boldsymbol{1}_{R}$ as a $p\times 1$ vector where we have ones in the places of restricted assets, and zeroes in the unrestricted assets. To continue with the example above $1_{R}=(1,0,0,0,1,0,1,0,0,0)^{\prime}$. Define $\bar{w}\mathrel{\mathop{:}}=\boldsymbol{1}_{R}^{\prime}\boldsymbol{w}-\boldsymbol{1}_{R}^{\prime}\boldsymbol{m}=\boldsymbol{1}_{R}^{\prime}\boldsymbol{w}_{d}$. Then $\boldsymbol{w}_{cp}$ is the difference between the portfolio and the benchmark portfolio when both benchmark and the portfolio is restricted. The constraints may take the form of certain stocks wanted to be excluded from the portfolio from the perspective of financial, moral, ethical and religious considerations. We allow for $r\to\infty$ as $p\to\infty$ as long as $10<\frac{r}{p}<$ 1. We allow for equality constraints of the form
$\bar{w}=\omega,$
where $\omega$ can be positive or negative. This representation is adopted by Bajeux-Besnainou et al. (2011). To give an example $\omega=0$ may show same restrictions applied to benchmark, $\boldsymbol{m}$, should apply to portfolio that we are forming, $\boldsymbol{w}$. To continue with that simple example with 10 assets above, if $\boldsymbol{1}_{R}^{\prime}\boldsymbol{w}=0.6$ means the weight of combined assets 1, 5 and 7 should be 0.6. If the weights are normalized to one and there is no short selling allowed, this means we form 60% of our portfolio from these three assets. These are very common restrictions in portfolio formation either by funds themselves or governments. This 60% can represent the percentage of stocks in a portfolio, for example. The optimization problem is:
@@
$\boldsymbol{w}_{cp}^{*}\mathrel{\mathop{:}}=\max_{\boldsymbol{w}_{d}\in R^{P}}[\boldsymbol{\mu}^{\prime}\boldsymbol{w}_{d}-\frac{\boldsymbol{\Xi}}{2}\boldsymbol{w}_{d}^{\prime}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{d}],\quad\mbox{subject to}\quad\boldsymbol{1}_{p}^{\prime}\boldsymbol{w}_{d}=0,\quad\boldsymbol{1}_{R}^{\prime}\boldsymbol{w}_{d}=\omega.$
@@
The solution is given by Proposition 1 in Bajeux-Besnainou et al. (2011). This is basically the solution to tracking error portfolio added to a fraction of the portfolio $\boldsymbol{l}$ where $\boldsymbol{l}$ can be thought as the weighted difference between restricted and unrestricted minimum variance portfolios.
$\boldsymbol{w}_{cp}^{*}=(\omega-\kappa w_{u})\boldsymbol{l}+\boldsymbol{w}_{d}^{*}=(\omega-\kappa w_{u})\boldsymbol{l}+(\boldsymbol{w}^{*}-\boldsymbol{m}),$ (7)
with
$\boldsymbol{l}\mathrel{\mathop{:}}=\frac{\boldsymbol{k}-\boldsymbol{a}}{w_{k}-w_{a}},$
and
$\boldsymbol{k}\mathrel{\mathop{:}}=\frac{\boldsymbol{\Theta}\boldsymbol{1}_{R}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}},\quad\boldsymbol{a}\mathrel{\mathop{:}}=\frac{\boldsymbol{\Theta}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}},$
@@
which are fully invested portfolios, and $\boldsymbol{a}$ is the minimum variance portfolio, and $\boldsymbol{k}$ can be thought roughly as its restricted counterpart. Note that $w_{k}$ is the addition of weights in the restricted minimum variance portfolio, and $w_{a}$ below represents the addition of weights in the minimum variance portfolio. Set
$w_{k}-w_{a}=\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}-\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}},$
@@
and $w_{k}\neq w_{a}$ since by assumption all assets in the portfolio cannot be restricted. We also have scalar
$w_{u}=\boldsymbol{1}_{R}^{\prime}[\frac{\boldsymbol{\Theta}\boldsymbol{\mu}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}}-\frac{\boldsymbol{\Theta}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}],$ (8)
@@
where $w_{u}$ can be thought as the subtraction of minimum variance portfolio weights from maximum Sharpe Ratio portfolio weights, when only restricted assets are taken into account. We can also rewrite the optimal joint restricted portfolio as
$\boldsymbol{w}_{cp^{*}}=\omega\boldsymbol{l}+\boldsymbol{w}_{d}^{*}-\kappa w_{u}\boldsymbol{l},$
where $\omega\boldsymbol{l}$ is the constrained minimum TE portfolio, and the additional weights $(w_{d}^{*}-\kappa w_{u}\boldsymbol{l})$ has zero weight on restricted assets, and provide additional return by increasing TE, as discussed by p.306 of Bajeux-Besnainou et al. (2011).

### v2

@@ 3 TE Constraint
with $\boldsymbol{m}:p\times 1$ benchmark portfolio of weights, $\boldsymbol{w}$ is the portfolio of weights that is tracking the benchmark portfolio $\boldsymbol{m}$, $\boldsymbol{\mu}:p\times 1$ vector of expected return on assets, and TE is the tracking error. Alternatively, we can write the optimization as
$\boldsymbol{w}_{d}^{*}:=\max_{\boldsymbol{w}_{d}\in R^{p}}[\boldsymbol{\mu}^{\prime}\boldsymbol{w}_{d}-\frac{\Xi}{2}\boldsymbol{w}_{d}^{\prime}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{d}],\quad\mbox{subject to}\quad\boldsymbol{1}_{p}^{\prime}\boldsymbol{w}_{d}=0,$
@@
with $\boldsymbol{w}_{d}=\boldsymbol{w}-\boldsymbol{m}$, $\infty>\Xi>0$, where $\Xi$ can be roughly seen as a risk aversion parameter. The solution to above problem is, $\boldsymbol{\Theta}:=\boldsymbol{\Sigma}_{y}^{-1}$, as in (1) of Bajeux-Besnainou et al. (2011)
@@
where $\frac{\boldsymbol{\Theta}\boldsymbol{\mu}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}}$ is the portfolio that maximizes Sharpe Ratio, and $\frac{\boldsymbol{\Theta}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}$ is the standard global minimum variance portfolio. Let $c,C$ be positive constants. Then, $\kappa$ represents the risk tolerance parameter, which is described in equation (5) of Bajeux-Besnainou et al. (2011), with a range of $0<c\leq|\kappa|\leq C<\infty$. For risk-averse investors $0<c\leq\kappa\leq C<\infty$. $\kappa,\boldsymbol{m}$ are given by investors and not estimated. The relation between TE, $\kappa,\Xi$ is given in Appendix A.3. The aim is to estimate optimal weights $\boldsymbol{w}_{d}^{*}$, and $\boldsymbol{w}^{*}:=\boldsymbol{w}_{d}^{*}+\boldsymbol{m}$. In that respect, we provide the following estimate $\hat{\boldsymbol{w}}_{d}=\kappa\left[\frac{\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}}-\frac{\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}\right],$ with $\hat{\boldsymbol{\mu}}=\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{y}_{t}$, with $\boldsymbol{y}_{t},$ $t=1,\cdots,T$, being the columns of $\boldsymbol{y}$ matrix. Note that since
@@
we have $\|\hat{\boldsymbol{w}}_{d}-\boldsymbol{w}_{d}^{*}\|_{1}=\|\hat{\boldsymbol{w}}-\boldsymbol{w}^{*}\|_{1}.$ We provide the following Theorem 1, which shows that portfolio weights can be consistently estimated in a large portfolio with TE constraint.
We also consider two other metrics: out-of-sample variance of a portfolio and out-of-sample Sharpe Ratio corresponding to $\hat{\boldsymbol{w}}$. Specifically, the out-of-sample variance estimate of the constrained portfolio is
@@
and the estimate of the out-of-sample Sharpe Ratio of the constrained portfolio is
@@
The Sharpe Ratio is $SR=\boldsymbol{w}^{*^{\prime}}\boldsymbol{\mu}/\sqrt{\boldsymbol{w}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}^{*}}$. Define the rate $r_{w1}:=\bar{s}l_{T}K^{3/2}$. We provide consistent estimation of portfolio weights and out-of-sample variance of the constrained portfolio.
Theorem 1
@@
$\|\hat{\boldsymbol{w}}-\boldsymbol{w}^{*}\|_{1}=\|\hat{\boldsymbol{w}}_{d}-\boldsymbol{w}_{d}^{*}\|_{1}=O_{p}(r_{w1})=O_{p}(\bar{s}l_{T}K^{3/2})=o_{p}(1).$
@@
$\left|\frac{\hat{\boldsymbol{w}}^{{}^{\prime}}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}}{\boldsymbol{w}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}^{*}}-1\right|=O_{p}(r_{w1}K^{5/2})=O_{p}(\bar{s}l_{T}K^{4})=o_{p}(1).$
@@
$\left|\left[\frac{\widehat{SR}}{SR}\right]^{2}-1\right|=O_{p}(r_{w1}K^{5/2})=O_{p}(\bar{s}l_{T}K^{4})=o_{p}(1).$
Remarks. 1. These are all new results in the literature and allow $p>T$ due to $l_{T}$ definition in (2), when both $p,T$ diverge. An illustrative example of a representative rate is given at the end of Section 2.
@@
3. If the precision matrix of errors is non-sparse, we can have $\bar{s}=p$. Then in this case $p<T$, we need $p=o(T^{1/4})$, by using $l_{T}$ definition in (2) for Theorem 1.
@@
4. We can also have an estimate for TE as $\widehat{TE}_{d}:=\sqrt{\hat{\boldsymbol{w}}_{d}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{d}}$, as an out-of-sample estimate. It can easily be shown that $|(\frac{\widehat{TE}_{d}}{TE})^{2}-1|$ will converge in probability to zero by following the proof of Theorem 1(ii), and the rate will be the same as in Theorem 1(ii). This estimator is consistent.
@@
5. Note that the variance and Sharpe Ratio have a slower rate of convergence compared to weight formation. This is clear by comparing Theorem 1(i) with Theorem 1(ii)-(iii).
6. In Appendix A.7, we analyze the differences analytically between the unconstrained portfolio weight and TE based portfolio weights in detail with proofs. In summary, estimating a TE constrained portfolio may be easier than the unconstrained portfolio in the finite samples due to the benchmark portfolio not being estimated if the risk tolerance parameters of both portfolios are the same. However, asymptotically, there is no difference; the rate of convergence of estimation errors to zero is the same and $r_{w1}$ in both cases, and the same analysis is true for variance and Sharpe Ratio estimation.
7. Our results use a diverging Sharpe Ratio of Maximum Sharpe Ratio (MSR) portfolio with a growing number of assets. This is a technical condition also used by Caner et al. (2023). If we impose a bounded Sharpe Ratio for MSR portfolio, $0<c\leq\boldsymbol{\mu}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}\leq C<\infty$, in Appendix B, Theorem B.1, Theorem B.3 clearly show we can have consistency but only with $p<T$ in two different cases with bounded Sharpe (MSR portfolio) ratio assumption.
8. Note that a possibility is to estimate the risk tolerance parameter $\kappa$ based on the lower bound for TE in Appendix A.3. We show that it is possible to estimate this parameter consistently in Appendix B.3. We provide a detailed proof. It can also be seen that since the rate matches Assumption 7(i), there will be no change in all the proofs here, as well as Theorems 2-3 below, if we use a constant $\kappa$.
4 Joint TE and Weight Constraints
@@
In this section we analyze weight constraints for a portfolio. Let $R$ represent the indices of restrictions and we define the cardinality of the set $R$ as $r:=|R|$. We allow for at least one restricted asset, and also, not all the assets can be restricted. In other words, the number of restrictions is $r$, $1\leq r<p$. To give an example, let $p=10$, and $R=\{1,5,7\}$, this means assets 1, 5, 7 are restricted in a 10 asset portfolio, and $r=3$. To make things a bit more clear in notation, define $\boldsymbol{1}_{R}$ as a $p\times 1$ vector where we have ones in the places of restricted assets and zeros in the unrestricted assets. To continue with the example above, $\boldsymbol{1}_{R}=(1,0,0,0,1,0,1,0,0,0)^{\prime}$. The benchmark portfolio is given, usually with weight restrictions. E.g., $\boldsymbol{m}=(0.1,0,0,0.1,0.2,0,0.3,0.1,0.1,0.1)$. Define $\omega:=\boldsymbol{1}_{R}^{\prime}\boldsymbol{w}-\boldsymbol{1}_{R}^{\prime}\boldsymbol{m}=\boldsymbol{1}_{R}^{\prime}\boldsymbol{w}_{d}$. Then $\omega$ is the difference between the portfolio and the benchmark portfolio when both the benchmark and the portfolio are restricted. The constraints may take the form of certain stocks being excluded from the portfolio from the perspective of regulatory, ethical and religious considerations. We allow for $r\to\infty$ as $p\to\infty$ as long as $0<\frac{r}{p}<1$ where $\omega$ can be positive or negative. This representation is adopted by Bajeux-Besnainou et al. (2011). To give an example, $\omega=0$ may show the same restrictions applied to the benchmark, $\boldsymbol{m}$, which should apply to a portfolio that we are forming, $\boldsymbol{w}$. To continue with that simple example with 10 assets above, $\omega=0$, which implies $\boldsymbol{1}_{R}^{\prime}\boldsymbol{w}=0.6$, the weight of combined assets 1, 5, and 7 should be 0.6. If the weights are normalized to one and there is no short selling allowed, this means we form 60% of our portfolio from these three assets. Inequality restrictions would be more common in portfolio formation either by funds themselves or regulators. We will discuss this in the next section. The optimization problem with equality constraint is:
@@
$\boldsymbol{w}_{cp}^{*}:=\max_{\boldsymbol{w}_{d}\in R^{P}}[\boldsymbol{\mu}^{\prime}\boldsymbol{w}_{d}-\frac{\boldsymbol{\Xi}}{2}\boldsymbol{w}_{d}^{\prime}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{d}],\quad\mbox{subject to}\quad\boldsymbol{1}_{p}^{\prime}\boldsymbol{w}_{d}=0,\quad\boldsymbol{1}_{R}^{\prime}\boldsymbol{w}_{d}=\omega.$
@@
The solution is given by Proposition 1 in Bajeux-Besnainou et al. (2011). This is basically the solution to TE portfolio added to a fraction of the portfolio $\boldsymbol{l}$ where $\boldsymbol{l}$ can be thought as the weighted difference between restricted and unrestricted minimum variance portfolios.
$\boldsymbol{w}_{cp}^{*}=(\omega-\kappa w_{u})\boldsymbol{l}+\boldsymbol{w}_{d}^{*},$ (7)
@@
with $\boldsymbol{l}:=\frac{\boldsymbol{k}-\boldsymbol{a}}{w_{k}-w_{a}},$ and $\boldsymbol{k}:=\frac{\boldsymbol{\Theta}\boldsymbol{1}_{R}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}},\,\boldsymbol{a}:=\frac{\boldsymbol{\Theta}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}},$ which are fully invested portfolios, and $\boldsymbol{a}$ is the global minimum variance portfolio, and $\boldsymbol{k}$ can roughly be considered as its restricted counterpart. See that $w_{k}:=\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}},\quad w_{a}:=\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}.$ Note that $w_{k}$ can be thought of the addition of weights in the restricted minimum variance portfolio, and $w_{a}$ represents the addition of weights in the minimum variance portfolio in a less restricted way.
@@
Note $w_{k}\neq w_{a}$ since, by assumption, not all assets in the portfolio are restricted ($r<p$). We also define
$w_{u}=\boldsymbol{1}_{R}^{\prime}\left[\frac{\boldsymbol{\Theta}\boldsymbol{\mu}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}}-\frac{\boldsymbol{\Theta}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}\right],$ (8)
@@
where $w_{u}$ can be thought as the subtraction of minimum variance portfolio weights from maximum Sharpe Ratio portfolio weights and that only restricted assets are taken into account.